Find the change in entropy when 3.0 kg of water is heated from 20°C to 80°C?

1
Expert's answer
2021-11-19T10:51:05-0500

The change in entropy

"\Delta S=\int_{T_1}^{T_2}\frac{dQ}{T}=\int_{T_1}^{T_2}\frac{cmdT}{T}"

"\Delta S=cm\ln\frac{T_2}{T_1}\\
=4200*3.0*\ln\frac{373}{293}=1321\:\rm J/ K"